A farmer's market has purchased 60 pumpkins for a total cost of $240.00. They expect that 20% of these pumpkins will remain unsold and that another eight will have to be sold at cost. If they wish to earn 30% gross profit on cost, what should be the selling price of the pumpkins?
Answer by ankor@dixie-net.com(12696)

A farmer's market has purchased 60 pumpkins for a total cost of $240.00. They expect that 20% of these pumpkins will remain unsold and that another eight will have to be sold at cost. If they wish to earn 30% gross profit on cost, what should be the selling price of the pumpkins?

60 - .20(60) = 48 pumpkins are sold
:
8 are sold at cost: 8*240/60 = $32
:
48 - 8 = 40 are sold at a profit
:
30% profit: 1.3*240 = $312 has to be obtained from the total sales
:
$32 received from the 8 pumpkins sold at cost>
:
312 - 32 = $280 has to be obtained from the remaining 40 pumpkins
:
280/40 = $7 for each pumpkin